OpenStudy (anonymous):

multiply the polynomial (a+5)3 the product is?

OpenStudy (anonymous):

\[(a+5)^{3}\]

OpenStudy (anonymous):

you want cube of a+5. if yes then simply apply formula on it.

OpenStudy (anonymous):

a+125 ?/idk im confused with this

OpenStudy (anonymous):

@uri this square formula not cube. @Avon yes you can calculate but you have to multiply one more time with a+5

OpenStudy (anonymous):

@uri im lost like what

OpenStudy (anonymous):

A+625 @gsRttN

OpenStudy (anonymous):

uri formula is right , after applying that formula again multiply the answer with a+5

OpenStudy (anonymous):

a+625 is ans ? i don't know answer i am not calculating it , i am just giving you solution , apply formula and try to calculate it own

OpenStudy (anonymous):

@gsRttN i dont even see @uri formla anymore

OpenStudy (anonymous):

#formula

OpenStudy (anonymous):

(a+b)^3 = a^3 + b^3 + 3ab(a+b) this is formula . try to apply it on your equation.

OpenStudy (anonymous):

\[a ^{4}+625\] @gsRttN

OpenStudy (anonymous):

a^4+625 is answer? @Avon

OpenStudy (anonymous):

@Avon you get it?

OpenStudy (anonymous):

no @gsRttN \[a ^{3}+15a ^{2}+75a+125 is the final answer\]

OpenStudy (anonymous):

yes according to your question.

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

@AkashdeepDeb @ivettef365 @PeterPan @Jhannybean @texaschic101

OpenStudy (anonymous):

\[(a+3)^{3}\]

OpenStudy (anonymous):

i dont get polynomials at all

OpenStudy (anonymous):

this is exactly same as last one. just replace 5 by 3. and you will get your answer

OpenStudy (akashdeepdeb):

Use this identity! \[(x+y)^3 = x^3 + y^3 + 3x^2y + 3xy^2\] Now solve for \[(a+3)^3 = a^3 + 3^3 + 3a^2.3 + 3.3^2.a\] Understood? :)

OpenStudy (anonymous):

thanks @ivettef365

OpenStudy (ivettef365):

yw :)

OpenStudy (texaschic101):

avon gave you the correct answer... a^3 + 15a^2 + 75a + 125

OpenStudy (anonymous):

thanks everyone for the help i appreciate it

OpenStudy (anonymous):

@Avon don't ever try to solve equation using any online calculator . try to solve it by hand. it will help you other wise it become habit and you will fail to solve such simple equation. so always try by own hand. Best of luck

OpenStudy (anonymous):

i will try @gsRttN